OpenStudy (anonymous):

Solve the following and Write Answer in Interval Notation. 2|4x-5|-4 ≤ 2x-2

OpenStudy (anonymous):

start with \[2|4x-5|\leq 2x+2\] then divide by 2 to get \[|4x-5|\leq x+1\] and then you have to work in cases

OpenStudy (anonymous):

that is, if \(x\geq\frac{5}{4}\) you know \(|4x-5|=4x-5\) so solve \[4x-5\leq x+1\] \[3x\leq 6\] \[x\leq 2\] so the interval \([\frac{5}{4},2]\) is good

OpenStudy (anonymous):

now repeat for \(x\leq \frac{5}{4}\) making \(|4x-5|=5-4x\) and solve \[5-4x\leq x+1\]

OpenStudy (anonymous):

what about the 4 in the beginning of the equation?

OpenStudy (anonymous):

i added 4 to both sides as a first step

OpenStudy (anonymous):

oh okay

OpenStudy (anonymous):

that why i wrote \(2|4x-5|\leq 2x+2\)

OpenStudy (anonymous):

when you solve the last inequality, you will get a contradiction check it and see (don't forget \(x<\frac{5}{4}\)

OpenStudy (anonymous):

so the "final answer: is just \([\frac{5}{4},2]\)

OpenStudy (anonymous):

alright thanks!

sam (.sam.):

I think its \[\frac{4}{5}<x<2\]

sam (.sam.):

2|4x-5|-4<2x-2 2|4x-5|<2x+2 |4x-5|<x+1 (4x-5)^2<x^2+2x+1 16x^2-40x+25<x^2+2x+1 15x^2-42x+24<0 (x-2)(5x-4)=0 x=2 x=4/5  Then 4/5<x<2